The unsigned binary equivalent of the decimal number 127 is O 01111111 11111111 10000000 00111111
Answer #1

The unsigned binary equivalent of the decimal number 127 is 01111111.

Here's the explanation:

when we are asked to find the binary equivalent of a decimal number,the most important is the modular division operation.

Modular division operation refers to the calculation of the remainder.

Given number is 127.

127%2=1 and 127/2=63

63%2=1 and 63/2=31

31%2=1 and 31/2=15

15%2=1 and 15/2=7

7%2=1 and 7/2=3

3%2=1 and 3/2=1

1%2=1 and 1/2=0

The operation is continued until quotient is 0.

Now take the remainder values from top to bottom.

Therefore the answer is 1111111.

The 8-bit representation of 127 is 01111111.

If you have any doubts regarding this topic, you can ask me.